# The gauge symmetry - unbroken and broken - in the QCD ground state

### v2

@@ Appendix A The full gluon propagator up to one loop
$N^{s}_{\rho\sigma}(l,q)$ $=$ $5l^{2}(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-20(ql)l_{\rho}l_{\sigma}+l^{2}(q^{2}-2(lq))\delta_{\rho\sigma}$ (A.5)
-- $2l^{2}q_{\rho}q_{\sigma}-10q^{2}l_{\rho}l_{\sigma}.$
Because of the two last terms and term $l^{2}q^{2}\delta_{\rho\sigma}$ in this equation, the subtracted integral (A.4) is only logarithmically divergent at the upper limit. So that the integral (A.3) is one of the tadpole-like terms which should be disregarded on the general basis due to the relations (6.6)-(6.7) within our approach. The subtracted integral (A.4) finally will contribute to the invariant function which appear in the full gluon propagator (7.5).
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lT^{0}_{\rho\sigma\alpha\beta}D_{\alpha\beta}(l)=\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (A.6)
Being the QD but already regularized constant, it is not connected to any external gluon momentum. In this respect it is different from all other tadpole-like terms which appear as a result of the subtraction scheme or in the NL iteration procedure. Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state 6 . We have shown that in the PT QCD it should be neglected on the general basis as well as all other tadpole-type terms. But in the QCD it should remain intact, while all other tadpole-types terms have to be removed from the theory, as it has been proven in section 6.
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l\ T^{0}_{\nu\zeta\mu}(l,p)D_{\nu\zeta}(l),$ (A.7)
where the gluon momentum $p$ is, in fact, zero, i.e., $p=0$ but it is convenient to go to this limit at the final step only. Both gluon propagators in the t’ Hooft-Feynman gauge are $D^{0}_{\mu\mu^{\prime}}(p)=\delta_{\mu\mu^{\prime}}/p^{2}$ and $D_{\nu\zeta}(l)=\delta_{\nu\zeta}d(l^{2})/l^{2}$, where $d(l^{2})$ is the corresponding invariant function of the full gluon propagator. Let us remind that such kind of the skeleton loop integrals are assumed to be regularized from above and below. Using now the Euclidean space Feynman rules for the corresponding vertices present in 6 , one arrives at
$T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ $=$ $-(2q+p)_{\mu^{\prime}}\delta_{\rho\sigma}+(q-p)_{\sigma}\delta_{\rho\mu^{\prime}}+(2p+q)_{\rho}\delta_{\sigma\mu^{\prime}},$
$T^{0}_{\nu\zeta\mu}(l,p)$ $=$ $2(l+p)_{\mu}\delta_{\nu\zeta}-l_{\zeta}\delta_{\nu\mu}-(2p+l)_{\nu}\delta_{\zeta\mu}.$ (A.8)
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p){2\over p^{2}}\int d^{4}l{d(l^{2})\over l^{2}}[3l_{\mu}+p_{\mu}].$ (A.9)
By the symmetry integration ($d^{4}l=l^{3}dl=(1/2)l^{2}dl^{2}$ and all the overall numerical numbers due to the integration over angular variables in four-dimensional Euclidean space 6 omitted) the first loop integral is zero, while the second one leads to
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p)p_{\mu}{2\over p^{2}}\times\tilde{\Delta}^{2}_{g}(d),$ (A.10)
where $\tilde{\Delta}^{2}_{g}(d)=\int d^{4}l(d(l^{2})/l^{2})$ is the QD but regularized skeleton loop integral, having the dimension of a mass squared. It is one of the constants present in the relations (6.7)-(6.8). It should be discarded on the general basis within our approach, i.e., put formally zero, so that $\tilde{\Delta}^{2}_{g}(d)=0$, even before the final $p^{2}=-M^{2}\rightarrow 0$ limit. Let us note that the cancellation of the pole $1/p^{2}=-1/M^{2}$ in eq. (B.10) can be always achieved by going to the dimensionless variable $x=l/M$ in the skeleton loop integral (A.9), i.e., this pole is not a problem here. Thus, $\Pi^{(4)}_{\rho\sigma}(q,p)=0$.
$\Pi^{(5)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l{l_{\mu}\over l^{2}}d_{gh}(l^{2}),$ (A.11)
where $d_{gh}(l^{2})$ is the invariant function for the full ghost propagator $G(l)$. The vertex $T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ is explicitly shown in the relations (A.8), and the remaining free gluon propagator is also the same. The altered rule 6 for the gluon-ghost vertex gives $(1/2)[(l_{\mu}+l{\mu})=l_{\mu}$ and the gluon momentum $p$ is zero again at the final stage. Since the nominator of this skeleton loop integral contains the loop variable linearly, then by the symmetry integration it is simply zero from the very beginning, i.e., $\Pi^{(5)}_{\rho\sigma}(q,p)=0$.

### v3

@@ Appendix B The origin of the conventional QCD
$\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)+\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (2.62)
The relation (3.13) in this case becomes
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.63)
As we already know, the quark contribution is always satisfied, i.e., $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0$, which leads to $\Delta^{2}_{q}=0$. Thus, the relation (B.3) is reduced to
$q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.64)
In complete analogy with the subtraction (4.19), let us now introduce the subtraction for the YM part as follows:
$\Pi^{s(YM)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\Pi^{YM}_{\rho\sigma}(0;D)$ (2.65)
$=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$
where $\Pi^{s(YM)}_{\rho\sigma}(0;D)=0$, by definition, and
$\Delta^{2}_{YM}(D)=\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (2.66)
as it follows from eq. (B.2).
Instead of the tensor decompositions (4.22), for the YM parts one gets
$\Pi^{s(YM)}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{s(YM)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(YM)}_{l}(q^{2}),$ (2.67)
and
$\Pi^{YM}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{YM}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{YM}_{l}(q^{2}).$ (2.68)
In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, all the invariant functions $\Pi^{s(YM)}_{t}(q^{2})$, $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtraction (B.5).
Substituting the relations (B.7)-(B.8) into the relations (B.4)-(B.5), by omitting the tedious tensor algebra derivations, one finally obtains
$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{YM}_{t}(q^{2})-{\Delta^{2}_{YM}(D)\over q^{2}},$
$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{\Delta^{2}_{YM}(D)\over q^{2}}.$ (2.69)
However, the invariant function $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the pole-type singularities, by definition, as underlined just above. So that one has to put the QD but regularized constant $\Delta^{2}_{YM}(D)=0$ on this mathematical ground. Taking into account that the quark constant $\Delta^{2}_{q}=0$ as well, one concludes that all the QD but regularized constants should be removed from theory, i.e., to be put zero as follows:
$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=\Delta^{2}_{t}(D)=0,$ (2.70)
because of the relation (B.6) and all of them are independent from each other. If $\Delta^{2}_{t}(D)=0$ then from the relation (B.2) it follows that $\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)$. But from the satisfied transverse relation $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and previous equality it follows that $q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)=0$ as well. Then from the relation (B.4) one gets $\xi_{0}=\xi$. The relations (B.10) and this equality determines the particular, or, equivalently, conventional QCD, investigated in section 6. Here it has been obtained by treating all the QD constants on the same footing, completely ignoring their principle difference from the physical point of view, which has been emphasized throughout this work.
This happened because in the previous derivations we deliberately have ignored that the subtraction (B.5), on account of the relation (B.2), in fact, looks like
$\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (2.71)
i.e., $\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{s(g)}_{\rho\sigma}(q;D)$, as it follows from eq. (4.18). Let us underline that the disappearance of the tadpole term is not a specific feature of our subtraction scheme. This always will happen in any subtraction scheme, since the constant minus the same constant is always zero, as pointed out in section 4. That is why explicitly present in the initial expressions it will remain in the final ones, compare eq. (2.5) with eq. (5.27), and thus with eq. (5.34).
Substituting further the tensor decomposition (B.7) and second of the decompositions (4.22) into the subtraction (B.11), and again omitting some tedious tensor algebra derivations, one obtains
$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},$
$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (2.72)
The general relation (B.4), on account of the first line of the subtraction (B.5), becomes
$q_{\rho}q_{\sigma}[\Pi^{s(YM)}_{\rho\sigma}(q;D)+\Pi^{YM}_{\rho\sigma}(0;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.73)
Substituting now the decomposition (B.7) and the relation $\Pi^{YM}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$, which follows from the relations (B.5), into the previous equation, and omitting again tensor algebra derivations, one obtains
$\Pi^{s(YM)}_{l}(q^{2})=-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{(\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D))\over q^{2}}.$ (2.74)
Equating now this relation to the second of the relations (B.12), and doing simple algebra, one finally arrives at
${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}=\Delta^{2}_{t}(D),$ (2.75)
which is nothing else but the general constraint (3.15). Note that in deriving this expression, we again use that from the satisfied transverse relation for the gluon part $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ it follows that $\Pi^{g}_{l}(q^{2})=0$. As in the case of the conventional QCD, the properties of the gluon degrees of freedom have been used at the last stage.
Even introducing the YM contribution to the full gluon self-energy in eq. (B.1), which hides the tadpole term, but correctly treated its subtraction (B.5) by the subtraction (B.11), one again arrives at the exact constraint (B.15), nevertheless. Treated it literally (i.e., by eq. (B.5) only) leads to the conventional QCD. In this case the significant piece of the information on the true dynamical and gauge structures of the QCD vacuum is losing. In other words, the conventional QCD hides the true role of the tadpole term. Contrary, in general QCD, based on the exact solution of eq. (B.15), this information is not losing and the true role of the tadpole has been revealed.
Concluding, let us emphasize that the exact constraint (3.15) has been obtained from the very beginning before introducing any proper subtraction scheme. This underlines its general and unique character. At the same time, the clarifying algebraic derivations performed here, leading to the relations (B.10), heavily depend on the not accurate (i.e., literally) treatment of the subtraction (B.5), as explained just above.
Appendix C Renormalized ST identity
Substituting eq. (3.16) into the ST identity (3.7), one obtains
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi=i{\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.76)
Let us introduce the non-perturbative (NP) renormalization constant for the tadpole term as follows: $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$, where $\Delta^{2}$ is finite and positive, by definition, while the NP renormalization constant $Z_{\Delta}$ depends on all the unphysical parameters, such as regulating parameters, mentioned above, etc. In principle, the string tension in the linear rising potential between heavy quarks has to be somehow related to the finite $\Delta^{2}$, since it is connected to the transverse projection operator structure, as underlined above. It is well-known that the ansatz $D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Delta^{2}/(q^{2})^{2}$ just leads to a such potential. After doing some algebra, the ST identity for the renormalized full gluon propagator becomes
$q_{\mu}q_{\nu}D^{R}_{\mu\nu}(q)=i\xi^{R}=i{\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}},$ (3.77)
where the renormalized gauge-fixing parameters are defined as follows: $\tilde{\xi}_{0}=Z_{\Delta}\xi_{0},\ \xi^{R}=Z_{\Delta}\xi$ and $D^{R}_{\mu\nu}(q)=Z_{\Delta}D_{\mu\nu}(q)$, i.e., the ST identity (C.1) is expressed in terms of the finite quantities only. The correct PT limit is maintained, of course.
Let us now demonstrate the interesting feature of the generalized gauge (C.2). It provides the self-consistency condition for the gauge choice in QCD as follows:
${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={aq^{2}\over q^{2}+a\Delta^{2}},$ (3.78)
[…]
${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={q^{2}\over\Delta^{2}},$ (3.79)
which is only satisfied at $q^{2}=0$, i.e., there is no any condition for the gauge-fixing parameter. The formulated self-consistency condition (C.3) points out on the inconsistency of the canonical gauge $\tilde{\xi}_{0}=\infty$ in QCD. Briefly, if everything is expressed in the terms of the finite quantities then the formal canonical gauge is forbidden to use in QCD, according to our approach to this theory.
Let us now apply the above-developed formalism to the renormalized massive gluon fields. For this purpose, it is instructive to re-write eq. (7.46) for the free massive gluon propagator in the terms of the renormalized quantities. But before it is convenient to introduce the following notations: $\tilde{\xi}_{0}=\tilde{\lambda}^{-1}$ and $M^{2}=M^{2}_{g}$, and the corresponding renormalization constant $Z_{\Delta}=1$ in this case. Then eq. (7.46) will look like
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{(1-\tilde{\lambda}^{-1})q_{\mu}q_{\nu}\over(q^{2}-\tilde{\lambda}^{-1}M^{2}_{g})}\right].$ (3.80)
From this observation it follows that our approach makes it possible for gluon to acquire mass dynamically without introducing into the theory some extra degrees of freedom such as Higgs fields [41], mentioned above, and $M^{2}_{g}$ can be treated as exactly defined gluon pole mass.
If one puts the expression (C.5) on the mass-shell $q^{2}=M^{2}_{g}$, then one arrives at
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{q_{\mu}q_{\nu}\over M^{2}_{g}}\right],$ (3.81)
which leads to the non-renormalizable behaviour in the PT $q^{2}\rightarrow\infty$ limit (i.e., it is constant and not $\sim 1/q^{2}$ as it is required for the renormalizable behaviour). However, the same expression (C.6) will be obtained if one expresses (C.5) in the canonical gauge ${\tilde{\lambda}}^{-1}=\infty$. As explained above, this gauge is not to be used in QCD. This means that within the generalized gauge-fixing function approach the free massive gluons cannot be put on the mass-shell. The finite values for the gauge-fixing parameter ${\tilde{\lambda}}^{-1}$ are only permitted in order to guarantee the theory to be renormalizable.The expression (C.5) has the correct PT limit, being off-mass-shell object at any finite ${\tilde{\lambda}}^{-1}$. So that, the canonical gauge is rather formal than real within the generalized gauge-fixing function approach to QCD. The canonical gauge and its connection to the structure of the full gluon propagator at the mass-shell will be investigated in our future work, where the NP renormalization theory of the massive gluon fields will be present in detail.
Appendix D Preservation of the gauge symmetry in the QED ground state
The propagation of the photon in the QED vacuum is formally described by eq. (2.1). Contrary to QCD (compare with eq. (2.2), the difference is that there is only one contribution to the full photon self-energy comes from the vacuum polarization tensor. Its skeleton term is similar present in Fig. 1 by the first diagram. Analytically it looks like
$\Pi^{ph}_{\rho\sigma}(q)\sim\int d^{4}l\gamma_{\rho}S(l-q)\Gamma_{\sigma}(l,q)S(l),$ (4.82)
where all the overall numerical factors are omitted, for simplicity. Here $S$ and $G$ represent the full electron propagator and the full electron-photon vertex, respectively. The full photon and its free counterpart are analytically present similar to eq. (3.8) and eq. (3.9), respectively, as well as their ST identities by relations (3.7) and (3.10).
Similar to the relation (3.11) for the transverse relation in this case, one obtains
$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (4.83)
It is well-known that because of the electric charge conversation, the transverse relation for the vacuum polarization tensor in QED is always satisfied, i.e., it is zero. Then from the previous equation immediately follows
$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)=0,\quad\rightarrow\quad\xi_{0}=\xi.$ (4.84)
Let us now introduce the corresponding subtraction scheme similar to the relation (4.19) and their independent tensor decompositions (4.22), namely
$\Pi^{s(ph)}_{\rho\sigma}(q)=\Pi^{ph}_{\rho\sigma}(q)-\Pi^{ph}_{\rho\sigma}(0)=\Pi^{ph}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{ph},$ (4.85)
where $\Pi^{s(ph)}_{\rho\sigma}(0)=0$, by definition, and
$\Pi^{s(ph)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(ph)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(ph)}_{l}(q^{2}),$
$\Pi^{ph}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{ph}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{ph}_{l}(q^{2}),$ (4.86)
respectively. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, the both invariant functions $\Pi^{s(ph)}_{t}(q^{2})$, $\Pi^{s(ph)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the initial subtraction (D.4).
Substituting all these decompositions (D.5) into the subtraction (D.4), and doing some tensor algebra derivations, one finally obtains
$\Pi^{s(ph)}_{t}(q^{2})$ $=$ $\Pi^{ph}_{t}(q^{2})-{\Delta^{2}_{ph}\over q^{2}},$
$\Pi^{s(ph)}_{l}(q^{2})$ $=$ $\Pi^{ph}_{l}(q^{2})+{\Delta^{2}_{ph}\over q^{2}}.$ (4.87)
From the first of the relations (D.3) and the second of the relations (D.5) it follows that $\Pi^{ph}_{l}(q^{2})=0$. Then the second of the relations (D.6) will be reduced to
$\Pi^{s(ph)}_{l}(q^{2})={\Delta^{2}_{ph}\over q^{2}},$ (4.88)
which is impossible since the corresponding invariant function can not have the pole-type singularities, by definition, as explained above. One has to put the QD but regularized constant to zero on this mathematical basis, i.e., $\Delta^{2}_{ph}=0$, and thus $\Pi^{s(ph)}_{l}(q^{2})=0$ as well.
As we already know, the both relations
$\xi=\xi_{0},\quad\Delta^{2}_{ph}=0$ (4.89)
are necessary conditions for the coincidence of the gauge symmetry of the QED Lagrangian with the gauge symmetry of its ground state (like it occurs for the conventional QCD). This is our proof of the Elitzur theorem [42] for the continuos abelian fields within our approach, based on the formalism of the dynamical equations of motion for the gauge particle Euclidean Green’s functions.